# XpookyNet: Advancement in Quantum System Analysis through Convolutional Neural Networks for Detection of Entanglement

### v3

@@ Abstract
The applications of machine learning models in quantum information theory have witnessed a surge in recent years, driven by the sophistication of quantum computing (QC). However, relying on inappropriate models has often resulted in suboptimal accuracy. A custom deep convolutional neural network (CNN), XpookyNet, has been developed to address this. XpookyNet is explicitly tailored for quantum systems analysis and aims to detect fundamental quantum attributes, such as entanglement in two- and three-qubit systems and potentially more qubit systems. It harnesses the power of CNNs to provide instantaneous and meticulous results. With an accuracy of 98.53% achieved in merely a few epochs, XpookyNet effectively handles the inherent complexity of quantum information, providing deeper insights into QC. The study also investigates quantum features and their relation to the purity of a density matrix directly related to noise in NISQ-era quantum computation. Preparing the density matrix in a format compatible with customary CNNs is vital in dissecting quantum systems. Quantum-applied CNNs must also be refined utilizing recent methods and extensive metrics to enhance and ensure their performance.
preprint: APS/123-QED
I Introduction
In quantum mechanics, an uncanny phenomenon known as quantum entanglement arises when two or more particles interact so that their quantum states become related in a way that the neither state $\psi_{AB}\neq\psi_{A}\otimes\psi_{b}$ nor $\langle A\otimes B\rangle\neq\langle A\rangle\otimes\langle B\rangle$ [1, 2]. Instantaneous changes in one particle are simultaneously mirrored in others, irrespective of their spatial separation, with the non-locality condition of $P(A=a,B=b)\neq P(A=a)\cdot P(B=b)$ holding [3]. Quantum Computing needs entanglement to form and rise, as it allows the system to perform multiple calculations simultaneously using many qubits, achieves secure and high key rates, and underpins quantum annealing and variational quantum algorithms to solve intractable problems and optimization problems [4, 5, 6, 7, 8].
@@ I Introduction
Quantifying and classifying entanglement beyond three qubits is an NP-Hard problem. The generalized concurrence still measures entanglement, but its calculation is computationally intensive for mixed states and it cannot pinpoint individual qubits [16]. Entanglement witnesses are another practical tool to detect entanglement in many-body quantum systems. A witness $W$ is a Hermitian operator that for all separable states, $\langle W\rangle_{\rho_{sep}}=\text{Tr}(W\rho_{sep})\geq 0$ but for some entangled states $\langle W\rangle_{\rho_{ent}}=\text{Tr}(W\rho_{ent})<0$ [17]. It detects entanglement without fully characterizing the system or performing a tomography; however, due to the exponential growth of variables with higher qubit counts, it requires optimization in high-dimensional spaces. Quantum witnesses can be optimized using machine learning (ML) because they can quickly identify patterns in large datasets, making them ideal for solving complex problems [18, 19]. Illustrations in Fig. 1 give a grasp of compartmentalizing methods for classifying states.
@@
Figure 1: Entangled or separable in Hilbert space. (a) Adjustment of a witness as a linear hyperplane and its optimization failure. (b) Entanglement witness optimization approaches, including linear: from $W_{1}$ to $W_{1}^{\prime}$, and nonlinear: from $W_{2}$ to $W_{2}^{\prime}$ [20, 21]. (c) Hindrance of precision due to the convexity of the target space. (d) Improvement of naive ML usage as convex witnesses isolator resulting in an ungeneralized model.
Among ML techniques, CNNs are versatile and robust, transforming research fields and impacting our daily lives [22]. They can attain accurate results on any dataset if it is properly encoded and the model is adapted to the data [23].
@@
A link between ML and quantum information (QI) has been extensively studied recently [24, 25, 26, 27, 28]. There has yet to be a model with rigorous accuracy for two-qubit entanglement detection problem, and the design of CNN models for QI applications remains relatively unexplored. The research needs to be more comprehensive in characterizing quantum system states beyond two qubits under the entanglement types. Additionally, the impact of noise on state identification regarding mixed states has yet to be extensively explored.
@@
Fig. 2 shows the research outline of this study. Initially, we create random density matrices and use them for training CNN. The density matrices produced for the dataset are categorized using the entanglement of formation (EoF). EoF underlies classifying the matrices based on their characteristics. The matrices are then pre-processed into an extended-tensor format to prepare them for subsequent processing, as explained in section II. In section III, we develop custom CNN models. The CNN model is designed and refined through a training and optimization process. Plateau machination, bracing layers, and branching features are added to enhance the model’s structure. The model is then subjected to testing to evaluate its performance. The resulting classification outcomes are used to assess the model’s applicability and make necessary adjustments. This approach ensures an accurately designed, trained, optimized, and tested CNN model. Section IV involves constructing confusion a matrix to assess the balance and symmetry of the results. It helps understand the model’s accuracy and errors made. Key metrics like ACC, MCC, and FNR are calculated to extensively analyze the study’s performance and reliability.
@@
II Quantum Data
There are two methods to combine quantum and classical computers for ML. The first method involves using a quantum computer to process and transform quantum data into a form that classical ML algorithms can understand [29]. The second method transfers data, like density matrix, from the quantum to the classical domain [30], which is being addressed in subsection . Both methods are in the early stages of research but have the potential to offer significant benefits.
II.1 Density Matrix and Entanglement
@@
$S(\rho_{A})=-\mathrm{Tr}\left[\rho_{A}\log_{2}{(\rho_{A})}\right],$ (3)
where $\rho_{A}$ represents the reduced density matrix of subsystem A.
II.2 Constructing Extended-Tensor
QI relies on complex numbers, but ordinary CNNs cannot accommodate them. We convert the density matrix into a three-dimensional tensor. Let $M$ be the initial $n\times n$ matrix constituted of complex numbers, and $M_{ij}=a_{ij}+\text{i}b_{ij}$ is defined as the complex number in the $i_{\text{th}}$ row and $j_{\text{th}}$ column of $M$. As illustrated in Fig. 3, we separate the density matrices into two matrices: one containing real numbers
$\sum_{j=0}^{n}\sum_{i=0}^{n}A_{ij}=\Re(M_{ij})$ (4)
and the other containing imaginary numbers
$\sum_{j=0}^{n}\sum_{i=0}^{n}B_{ij}=-\Im(M_{ij})*\text{i}$ (5)
Finally, we obtain a tensor that forms the input data as follows:
$\sum_{k=0}^{1}\begin{cases}C_{ijk}=A_{ij},&k=0\\ C_{ijk}=B_{ij},&k=1\end{cases}$ (6)
Tensor $C$ congregates matrix A and B, which have separated the real and imaginary parts into different layers of the 3D on each row $i$ and column $j$.
Figure 3: The process of converting complex number data to real number data.
II.3 Two-qubit data generation
To generate the density matrix for a two-qubit system, we generate a random complex matrix $M$ of size $4\times 4$ with elements of $M_{i,j}=a+ib$ where $a$ and $b$ are drawn from a Gaussian distribution. A Hermitian matrix $H$ is formed from the random complex matrix $M$ by calculating $H=M+M^{\dagger}$ to ensure that $H$ is Hermitian. Then $\text{Tr}(H^{2})>0$ is checked to ensure $\langle\psi|\rho|\psi\rangle\geq 0$. Then the density matrix $\rho$ is formed from the Hermitian matrix $H$ by normalizing it, i.e., $\rho=H/\text{Tr}(H)$, to ensure that the $\text{Tr}(\rho)=1$. QuTiP facilitates this procedure [33]. We regulate our density matrices to have 75% non-zero elements so it is more likely to generate a complex, mixed quantum state and feature significant quantum superposition. However, since begetting random two-entangled $\rho$ is approximately three times more likely than begetting separable $\rho$, it is vital to balance our dataset to prevent bias and improve CNN’s accuracy.
We generate one million $4\times 4$ random density matrices as our dataset [34]. This dataset contains 500,000 entangled and 500,000 separable $\rho$ that are explained in Appendix.A. The matrices are labeled by calculating the EoF of each matrix using the Qiskit [35].
$\text{Label}(\rho)=\begin{cases}\text{Ent},&\text{EoF}(\rho)>0\\ \text{Sep},&\text{EoF}(\rho)=0\end{cases}$ (7)
@@ II.3 Two-qubit data generation
II.4 Three-qubit data generation
@@
In the three-qubit quantum system Hilbert space onwards, it is not pervasive to only divide states into entangled or separate. As depicted in Fig. 4(a), entangled states itself are divided into four types $|\Psi_{ABC}\rangle$, $|\Psi_{AB|C}\rangle$, $|\Psi_{BC|A}\rangle$, and $|\Psi_{AC|B}\rangle$. We detailed in Appendix.B how to obtain each type but superficially saying, partial entanglements $|\Psi_{AB|C}\rangle$, and $|\Psi_{BC|A}\rangle$ are provided with a $\otimes$ operator between an entangled $|\Psi_{i,j}\rangle$ and $|\Psi_{k}\rangle$ but $|\Psi_{AC|B}\rangle$ is not that easy. Also, $|\Psi_{ABC}\rangle$ are demanding, which puts our focus on known three-qubit states GHZ, W state, and Graph state. Fig. 4(b) physically demonstrates the entangled states’ type, which exerts randomness by single random unitary operations $U_{r}^{A,B,C}$ and entanglement by $U^{Q}_{Ent.}$ operator, where $Q$ determines monolith entangled qubits. The number of $Q$ over $n$ qubits is as follows:
@@
We create 250,000 density matrices for three-qubit states, divided into five balanced categories as a dataset, which takes only 102 seconds to prepare 100% pure states when running on a regular CPU.
III Designing a CNN
@@
Figure 5: XpookyNet’s whole scheme. Blue layers show the principal artery and green layers show the branches. The last layer consists of two dense layers.
@@
To understand CNN performance, a loss function is needed. The loss function categorical cross entropy (CCE) is calculated for a given number of classes, denoted as $C$.
@@
In CCE’s Equation (11), $y_{i}$ represents the actual label for class $i$, and $\hat{y_{i}}$ denotes the predicted SoftMax probability for class $i$. Considering only entangled or not, the proposed equation (2) simplifies to a binary classification problem by setting $C=2$. The modified equation becomes:
@@
The back-propagation algorithm optimizes weight and bias in equation (9) over loss functions:
$\delta_{j}^{k}\equiv\frac{\partial a_{j}^{k}}{\partial E}$ (14)
Here, $\delta_{j}^{k}$ is the loss at layer $k$ for node $j$, $a_{j}^{k}$ is the activation of node $j$ at layer $k$, and $E$ is the loss function. The algorithm is used to update the weights of the network during training. The loss $\delta_{j}^{k}$ at layer $k$ is dependent on the loss $\delta_{k}^{k+1}$ at the next layer $k+1$. Thus, loss flows backward from the last layer to the first layer.
@@ III.1 Proposed CNN
The proposed CNN, XpookyNet, can determine the spooky action at a distance, its amount, and the entangled qubits by using varied activation functions in the last layer accordingly.
XpookyNet has parallel layers with expanded kernel sizes, and it uses conventional convolutional layers that can be replaced with more efficient ”separable convolutional layers.” Batch Normalization (BN) stabilizes CNNs and improves their performance, accuracy, and speed [38]. Separable convolutions make CNN training easier, faster, and more accurate while reducing computational complexity and overfitting. However, they may be less effective with low-dimension tensors and require more layers to achieve accuracy, similar to traditional CNNs [39].
III.2 Training CNNs
XpookyNet uses the stochastic gradient descent (SGD) optimizer with momentum and saves the most accurate model with minimum loss. The learning rate is adjusted when the loss reaches a plateau [40]. Parallel paths help prevent vanishing gradients again by giving a shortcut to backward flow [41]. Also, balancing and shuffling the data improves results.

### v4

@@ Abstract
Quantum system attributes, notably entanglement, are indispensable in manipulating quantum information tasks. Ergo, machine learning applications to help harness the sophistication of quantum information theory have surged. However, relying on unsuited prototypes and not bridling quantum information data for usual processors has often resulted in sub-optimal efficaciousness and inconvenience. We develop a custom deep convolutional neural network, XpookyNet, which is streamlined with respect to the interrelationships of density matrices of two-qubit systems to underpin three-qubit systems, breaking new ground to more qubit systems by their subsystems. Comparative implementation of XpookyNet provides instantaneous and meticulous results with an accuracy of 98.53% in merely a few epochs. XpookyNet effectively handles the inherent complexity of quantum information, equipping deeper insights into many-body systems as a bedrock. The study also investigates quantum features and their relation to the purity of a density matrix directly related to noise and system decoherence in NISQ-era quantum computation. Preparing the density matrix in a compact and compatible format with customary convolutional neural networks plays a determining role in dissecting quantum features. The procedure renders a convex criterion that detects entanglement and is a yardstick for quantum system coherence.
keywords:
Quantum Entanglement Detection, Machine Learning, Deep Learning, Convolutional Neural Network, Partial Entanglement, Quantum Data.
1 Introduction
In quantum mechanics, an uncanny phenomenon known as quantum entanglement arises when two or more particles interact so that their quantum states become related in a way that the neither state $|\psi_{AB}\rangle\neq|\psi_{A}\rangle\otimes|\psi_{B}\rangle$ nor $\langle A\otimes B\rangle\neq\langle A\rangle\otimes\langle B\rangle$ [1, 2]. Instantaneous changes in one particle are simultaneously mirrored in others, irrespective of their spatial separation, with the non-locality condition of $P(A=a,B=b)\neq P(A=a)\cdot P(B=b)$ holding [3]. Quantum computing needs entanglement to form and rise, as it allows the system to perform multiple calculations simultaneously using many qubits, achieves secure and high key rates, and underpins quantum annealing and variational quantum algorithms to solve intractable problems and optimization problems [4, 5, 6, 7, 8].
@@ 1 Introduction
Quantifying and classifying entanglement beyond three qubits is an NP-Hard problem. The generalized concurrence still measures entanglement, but its calculation is computationally intensive for mixed states and it cannot pinpoint individual qubits [16]. Entanglement witnesses are another practical tool to detect entanglement in many-body quantum systems. A witness $\mathcal{W}$ is a Hermitian operator that for all separable states, $\langle\mathcal{W}_{\rho_{sep}}\rangle=\text{Tr}(\mathcal{W}\rho_{sep})\geq 0$ but for some entangled states $\langle\mathcal{W}_{\rho_{ent}}\rangle=\text{Tr}(\mathcal{W}\rho_{ent})<0$ [17]. It detects entanglement without fully characterizing the system or performing a tomography; however, due to the exponential growth of variables with higher qubit counts, it requires optimization in high-dimensional spaces. Quantum witnesses can be optimized using machine learning (ML) because they can quickly identify patterns in large datasets, making them ideal for solving complex problems [18, 19]. Illustrations in Fig. 1 give a grasp of compartmentalizing methods for classifying states [20, 21].
@@
Figure 1: Entangled or separable in Hilbert space. (a) Adjustment of a witness as a linear hyperplane and its optimization failure. (b) Entanglement witness optimization approaches, including linear: from $\mathcal{W}_{1}$ to $\mathcal{W}_{1}^{\prime}$, and nonlinear: from $\mathcal{W}_{2}$ to $\mathcal{W}_{2}^{\prime}$. (c) Hindrances to precision due to the convexity of the target space. (d) Improvement of naive ML usage as a convex witness isolator resulting in an ungeneralized model.
Convolutional neural networks (CNNs) have revolutionized computer vision. They exploit hierarchical layers of convolution operations to exploit structured patterns in 3D tensors like images. To detect more complex structures, CNNs use convolutional layers as filters to identify edges, textures, and other details. By using hierarchical approaches, CNNs are able to understand spatial hierarchies and correlations within structured data [22].
Density matrices of quantum states share a structured nature analogous to the grid structure of images. They epitomize the coherence of different quantum states, similar to pixels in an image, arranged in a grid-like fashion. The application of CNNs to density matrices enables the identification of critical quantum properties and uncovers patterns that might be difficult to discern with traditional analyses [23].
@@
The link between ML and quantum information (QI) has recently been extensively studied [24, 25, 26]. There has yet to be a model with sufficiently rigorous accuracy for the two-qubit entanglement detection problem, and the design of CNN models for QI applications remains relatively unexplored. The research must characterize quantum system states more comprehensively beyond two qubits under the entanglement categories.
@@
Figure 2 shows the research outline of this study. Initially, we create random density matrices and use them for training CNN. The density matrices produced for the data set are labeled using entanglement of formation (EoF). The EoF underlies the classification of the matrices on the basis of their characteristics. The matrices are then pre-processed into an extended-tensor format to prepare them for subsequent processing, as explained in section 2. In section 3, we develop custom CNN models. The CNN model is designed and refined through a training and optimization process. Plateau machination, data normalization, and branching features are added to enhance the model structure. The model is then subjected to testing to evaluate its performance. The resulting classification outcomes are used to assess the model’s applicability and make necessary adjustments. This approach ensures an accurately designed, trained, optimized, and tested CNN model. Section 4 involves constructing a confusion matrix to assess the balance and symmetry of the results. It helps to understand the accuracy of the model and the errors made. Key metrics like ACC, MCC, and FNR are calculated to analyze the study’s performance and reliability extensively.
@@
2 Quantum Data
There are two methods to combine quantum and classical computers for ML. The first method involves using a quantum computer to process and transform quantum data into a form that classical ML algorithms can understand [27]. The second method transfers data, like density matrix, from the quantum to the classical domain [28], which is being addressed in the following subsection. Both methods are in the early stages of research but have the potential to offer significant benefits.
2.1 Density Matrix and Entanglement
@@
$S(\rho^{A})=-\mathrm{Tr}\left[\rho^{A}\log_{2}{(\rho^{A})}\right],$ (3)
where $\rho^{A}$ represents the partial trace of density matrix $\rho$ regarding the other subsystem i.e., subsystem $B$.
2.2 Constructing Extended-Tensor
QI relies on complex numbers, but ordinary CNNs cannot accommodate them. We convert the density matrix into a three-dimensional tensor, which is less voluminous than regular storage of the complex number $a+\text{i}\,b$ to the $2\times 2$ matrix of:
$\begin{bmatrix}a&-b\\ b&a\end{bmatrix}.$
Let $\mathbf{M}_{n\times n}$ be the initial matrix made up of complex numbers and $M_{ij}=a_{ij}+\text{i}\,b_{ij}$ be defined as the complex number in the row $i^{\text{th}}$ and the column $j^{\text{th}}$ of $\mathbf{M}_{n\times n}$. As illustrated in Fig. 3, we separate the density matrices into two matrices $\mathbf{A}_{n\times n}$ and $\mathbf{B}_{n\times n}$. Tensor $\mathbf{A}_{n\times n}$ contains real part coefficient,
$A_{ij}=\Re(M_{ij})=a_{ij}\ \ \ \text{for}\ \ \ i,j=1,2,...,n$ (4)
and tensor $\mathbf{B}_{n\times n}$ contains imaginary part coefficient,
$B_{ij}=\Im(M_{ij})=b_{ij}\ \ \ \text{for}\ \ \ i,j=1,2,...,n$ (5)
Finally, we obtain a 3D order tensor that forms the input data as follows:
$\begin{cases}C_{ij0}=A_{ij}&\ \ \ \text{for}\ \ \ i,j=1,2,...,n\\ C_{ij1}=B_{ij}&\ \ \ \text{for}\ \ \ i,j=1,2,...,n\end{cases}$ (6)
Tensor $\mathbf{C}_{n\times n\times 2}$ congregates matrix $\mathbf{A}_{n\times n}$ and $\mathbf{B}_{n\times n}$, which have separated the real and imaginary parts into different layers of the 3D on each row $i$ and column $j$.
Figure 3: The process of converting complex number data to real number data (extended data).
2.3 Two-qubit data generation
To generate the density matrix for a two-qubit system, we generate a random complex matrix $\mathbf{M}_{4\times 4}$ with elements of $M_{i,j}=a_{ij}+\text{i}\,b_{ij}$ where $a_{ij}$ and $b_{ij}$ are drawn from a Gaussian distribution. A Hermitian matrix $H$ is formed from the random complex matrix $\mathbf{M}_{4\times 4}$ by calculating $H=\mathbf{M}+\mathbf{M}^{\dagger}$ to ensure that $H$ is Hermitian. Then $\text{Tr}(H^{2})>0$ is checked to ensure $\langle\psi|\rho|\psi\rangle\geq 0$. Then the density matrix $\rho$ is formed from the Hermitian matrix $H$ by normalizing it, i.e., $\rho=H/\text{Tr}(H)$, to ensure that the $\text{Tr}(\rho)=1$. QuTiP facilitates this procedure [31]. We regulate our density matrices to have 75% non-zero elements so it is more likely to generate a complex, mixed quantum state and feature significant quantum superposition. However, since begetting random two-entangled $\rho$ is approximately three times more likely than begetting separable $\rho$, it is vital to balance our dataset to prevent bias and improve CNN’s accuracy.
We generate one million random density matrices as our dataset [32]. This dataset contains 500,000 entangled and 500,000 separable $\rho$ that are explained in Appendix A . The matrices are labeled by calculating the EoF of each matrix using the Qiskit [33].
$\text{Label}(\rho)=\begin{cases}\text{Ent.},&\text{EoF}(\rho)>0\\ \text{Sep},&\text{EoF}(\rho)=0\end{cases}$ (7)
@@ 2.3 Two-qubit data generation
2.4 Three-qubit data generation
@@
In the three-qubit quantum system Hilbert space onwards, it is not pervasive to only divide states into entangled or separate. As depicted in Fig. 4(a), entangled states itself are divided into four types $|\Psi_{ABC}\rangle$, $|\Psi_{AB|C}\rangle$, $|\Psi_{BC|A}\rangle$, and $|\Psi_{AC|B}\rangle$. We detailed in Appendix B how to obtain each type but superficially saying, partial entanglements $|\Psi_{AB|C}\rangle$, and $|\Psi_{BC|A}\rangle$ are provided with a tensor operator ($\otimes$) between an entangled $|\Psi_{i,j}\rangle$ and $|\Psi_{k}\rangle$ but $|\Psi_{AC|B}\rangle$ is not that easy. Also, $|\Psi_{ABC}\rangle$ are demanding, which puts our focus on known three-qubit states: GHZ state, W state, and Graph state. Fig. 4(b) physically demonstrates the entangled states’ type, which exerts randomness by single random unitary operations $U_{Ran.}^{A,B,C}$ and entanglement by $U^{Q}_{Ent.}$ operator, where $Q$ determines monolith entangled qubits. The number of $Q$ over $N$ qubits is as follows:
@@
We create 250,000 density matrices for three-qubit states, divided into five balanced categories as a dataset, which takes only 102 seconds to prepare states with purity of 100% when running on a regular CPU.
3 Designing a CNN
@@
To understand CNN performance, a loss function is needed. In a classification task, the loss function categorical cross entropy (CCE) is calculated for a given number of classes, denoted as $C$.
@@
In CCE’s Equation (11), $y_{i}$ represents the actual label for class $i$, and $\hat{y_{i}}$ denotes the predicted SoftMax probability for class $i$. In 2-qubit system, considering only entangled or not, the proposed equation (2) simplifies to a binary classification problem by setting $C=2$. The modified equation becomes:
@@
Furthermore, the model functions for regression task to estimate the EoF value of each density matrix can be define by shifting the loss function to mean square error (MSE), which is obtained as follows:
$L_{MSE}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}(y_{i}-\hat{y}_{i})^{2}$ (14)
where $N_{s}$ is the total number of observations, $y_{i}$ is the EoF of the $i^{th}$ instance, and $\hat{y}_{i}$ is the corresponding model prediction.
The back-propagation algorithm optimizes weight and bias in equation (9) over loss functions $L$ as follows:
$\theta^{k-1}=\theta^{k}-\alpha\frac{\partial L}{\partial\theta^{k}}$ (15)
Here, $\theta^{k}$ represents the parameters at layer $k$.
@@ 3.1 Proposed CNN
The phenomenon of entanglement, often referred to as ”spooky action at a distance”, is central to this study, which introduces a CNN framework named XpookyNet. This framework is designed to identify entangled states and quantify their degree by employing various optimization and learning techniques. Figure 5 depicts the entire XpookyNet architecture, which incorporates several enhancement methods. The core structure of XpookyNet comprises 10 convolutional layers (Simple Conv.) and two fully connected neural networks (NN). The enhancement methods of XpookyNet include (1) batch normalization combined with separable convolutional layers (B.N. Sep.), (2) three-branch convolutional layers (Brch.) and (3) the learning rate reduction method on plateaus (Plat.).
In the context of batch normalization with separable convolutional layers, batch normalization improves the stability, performance, accuracy, and speed of the CNN [36]. In addition, separable convolutional layers make CNN training more efficient, faster, and precise by lowering computational complexity and mitigating overfitting. However, they might be less effective with low-dimensional tensors and typically necessitate more layers to achieve the desired accuracy [37].
Figure 5: XpookyNet’s whole scheme. Blue layers show the principal artery and green layers show the branches. The last layer consists of two fully connected dense layers.
3.2 Training CNNs
XpookyNet uses the stochastic gradient descent (SGD) optimizer with momentum and saves the most accurate model with minimum loss. The learning rate is adjusted when the loss reaches a plateau [38]. Parallel paths help prevent vanishing gradients as well by giving a shortcut to backward flow [39]. Also, balancing and shuffling the data improves results.